use the variable expression 3x-1.
evaluate the expression when x=-3,-2,-1,0,1,2,and 3.

Answer by elima(1433)

3x-1.
evaluate the expression when x=-3,-2,-1,0,1,2,and 3
Substitute each number for the x;
3(-3)-1 = -9-1=-10
3(-2)-1 = -6-1=-7
... and so on until you have done each number.